OpenStudy (anonymous):

Help! if f(x)= 6x and g(x)= 4x+1 24x+6 24x^2+6x 24x+1 10x+1 I picked the first one.

OpenStudy (anonymous):

what are you trying to find?

OpenStudy (anonymous):

The fog(x)

OpenStudy (anonymous):

so when you know the functions, you just have to put them in the correct spot. in \((f \circ g)(x)\) what is the inside function, f or g?

OpenStudy (anonymous):

F, right?

OpenStudy (anonymous):

yep, so it goes inside f, like this:\[\left( f \circ g \right)\left( x \right)=f \left( g \left( x \right) \right)=f \left( 4x+1 \right)=6\left( 4x+1 \right)\]

OpenStudy (anonymous):

Wait, so I was right?

OpenStudy (triciaal):

inside function would be g

OpenStudy (anonymous):

yes, but you need to be confident and sure. when you are, then will explain how you got what you got and the priniclple you followed to prove that you are correct. you're certainly getting there, yes?

OpenStudy (anonymous):

Yes, I get it way more now actually.

OpenStudy (anonymous):

awesome!